# RealTraj: Towards Real-World Pedestrian Trajectory Forecasting

### v2

@@ 4.1 Datasets
Joint Track Auto (JTA) Dataset: The JTA dataset [15] is a large-scale synthetic dataset comprising $256$ training sequences, $128$ validation sequences, and $128$ test sequences set in urban scenarios, generated using a video game engine.
ETH-UCY: The ETH [45] and UCY [33] datasets serve as established benchmarks for human trajectory forecasting, encompassing five scenes. We utilize the leave-one-out validation strategy as described in prior works [24, 37].
Stanford Drone Dataset (SDD): The SDD [48] dataset features over $11,000$ unique pedestrians across $20$ scenes, captured from a bird’s-eye view using a drone. Following the previous trajectory forecasting methods [37], we use the standard setup and train-test split.
TrajImpute: The TrajImpute [9] dataset simulates missing values within ETH-UCY, offering a framework for evaluating trajectory imputation and prediction models.
Following established conventions, we predict $12$ future timesteps based on $9$ past timesteps for the JRDB and JTA datasets, and $12$ future timesteps based on $8$ past timesteps for the ETH-UCY, SDD, and TrajImpute datasets, all at a frame rate of $2.5$ fps.
@@ 4.2 Evaluation Metrics
We adopt the widely-used evaluation metrics Average Displacement Error (ADE) and Final Displacement Error (FDE). ADE calculates the average error across all ground truth and predicted positions in the trajectory, while FDE measures the displacement between the endpoints of the ground truth and predicted trajectories.
@@ 4.3 Implementation Details
We use the ORCA crowd simulator [58], implemented by [47], to generate synthetic trajectories. Up to $40$ pedestrians are placed in a $15\text{m}\times 15\text{m}$ environment with up to $20$ static primitive obstacles. Examples of generated synthetic trajectories are shown in Fig. 3. We generate $2,000$ sequences for pretraining. We train our model using the Adam optimizer, starting with a learning rate of $1\times 10^{-4}$, which is reduced by a factor of $0.1$ after completing $80\%$ of the total $200$ epochs for pretraining and $20$ epochs for fine-tuning. During the first $100$ epochs of pretraining, the loss weights $\alpha$, $\beta$, and $\gamma$ are set to $1$, 00, and 00, respectively; for the final $100$ epochs, these weights are set to 00, $100$, and $0.1$. A corruption ratio of $30\%$ is applied. Hyperparameters were selected through a coarse-to-fine grid search and step-by-step tuning. The model is trained with a batch size of $16$ on a single NVIDIA TITAN RTX GPU.
Table 2: Comparison with state-of-the-art methods on the JTA, JRDB, ETH-UCY, and SDD. The unit for ADE and FDE is meters for JRDB, JTA, and ETH-UCY, and pixels for SDD. The best results are bolded, and the second-best results are underlined.
Setting Method Input JRDB JTA ETH-UCY SDD
ADE FDE ADE FDE ADE FDE ADE FDE
Supervised Vanilla-LSTM [1]${}_{\text{CVPR'}16}$ Trajectory 0.42 0.83 1.44 3.25 - - - -
Occupancy-LSTM [1]${}_{\text{CVPR'}16}$ 0.43 0.85 1.41 3.15 - - - -
Social-LSTM [1]${}_{\text{CVPR'}16}$ 0.47 0.95 1.21 2.54 0.72 1.54 - -
Social-GAN-ind [24]${}_{\text{CVPR'}18}$ 0.50 0.99 1.66 3.76 0.74 1.54 - -
Transformer [22]${}_{\text{ICPR'}21}$ 0.56 1.10 1.56 3.54 0.54 1.17 - -
Directional-LSTM [32]${}_{\text{T-ITS'}22}$ 0.45 0.87 1.37 3.06 - - - -
Dir-social-LSTM [32]${}_{\text{T-ITS'}22}$ 0.48 0.95 1.23 2.59 - - - -
Trajectron++ [52]${}_{\text{ECCV'}20}$ 0.40 0.78 1.18 2.53 0.53 1.11 - -
Autobots [21]${}_{\text{ICLR'}22}$ 0.39 0.80 1.20 2.70 0.52 1.05 - -
TUTR [53]${}_{\text{ECCV'}20}$ 0.53 1.12 17.38 35.04
EqMotion [65]${}_{\text{CVPR'}23}$ 0.42 0.78 1.13 2.39 0.49 1.03 16.75 33.72
Social-Trans [50]${}_{\text{ICLR'}24}$ 0.37 0.72 0.92 1.84 0.51 1.03 17.96 38.66
RealTraj (Ours) Detection 0.35 0.67 0.92 1.84 0.51 1.03 15.99 31.83
Weakly-Supervised RealTraj (Ours) Detection 0.36 0.71 1.04 2.11 0.52 1.07 16.02 31.98
Refer to caption (b) Miss-Detection Refer to caption (a) Error Localization Refer to caption (c) Identity Switch Refer to caption (d) Combined Error
@@ 4.4 Few-shot Evaluation
We first demonstrate the effectiveness of RealTraj with limited real-world labeled data. We randomly selected subsets of $0.1\%$, $0.5\%$, $1\%$, $2\%$, and $5\%$ of the dataset, fine-tuned the model on each subset, and evaluated it on the test set on JRDB, JTA, ETH-UCY, and SDD. Since no prior work focuses on a Weakly-Supervised trajectory forecasting method trained with weak detection annotations (denoted as “Weakly-Supervised”), we compare our model with supervised approaches trained with trajectory annotations (denoted as “Supervised”). Specifically, we compare against the current state-of-the-art supervised methods, EqMotion [65] and Social-Transmotion (Social-Trans) [50].
As shown in Tab. 1, our method consistently outperforms these prior methods across all data regimes on JRDB. In the most challenging scenario with only $0.1\%$ of annotations, RealTraj achieves improvements of $4.4\%$, $20\%$, and $4.8\%$ in terms of ADE on JRDB, ETH-UCY, and SDD, respectively, compared to the best-performing previous supervised model, EqMotion. In other data regimes, RealTraj achieves results comparable to supervised methods despite requiring only weak detection annotations during fine-tuning. We also provide results for RealTraj trained in a supervised manner with trajectory annotations (Supervised-RealTraj), which substantially improves over prior supervised methods. Notably, RealTraj closely matches the performance of Supervised-RealTraj on JRDB and SDD across all data regimes. These findings confirm that the proposed model significantly improves performance with limited and weak detection annotations, reducing both data collection and annotation costs, mentioned in Sec. 1.
@@ 4.5 Comparison with State-of-the-art
We further compare RealTraj with a wide range of current supervised approaches under a standard $100\%$ labeled evaluation setting, depicted in Tab. 2. The results demonstrate that our framework outperforms all other methods on the JRDB, JTA, and SDD datasets in terms of ADE, surpassing existing approaches by over $4.4\%$ in ADE and $5.2\%$ in FDE on the SDD dataset. Additionally, Supervised-RealTraj achieves the best performance among all methods on JRDB, JTA, ETH-UCY, and SDD in terms of ADE.
4.6 Robustness Evaluation
The robustness of the proposed method against perception errors (detection and tracking errors) is compared with that of the Social-Trans [50] and EqMotion [65], which are the best-performing state-of-the-art method considering both ADE and FDE metrics, as shown in  Tab. 2. We synthetically generated two types of detection errors—miss-detections and localization errors—along with one type of tracking error, specifically identity switches, and combined these errors for evaluation. Miss-detections were simulated by setting past coordinates to zero, while localization errors were introduced by adding Gaussian noise to past coordinates. Identity switches were simulated by swapping identities with nearby pedestrians within a 5-meter radius. Both detection and tracking errors were introduced simultaneously for combined errors. To enable comparison methods to handle missing values in past trajectories, linear interpolation, and extrapolation were applied to fill in missing values before feeding the data into the model.
Table 3: Comparison with the current state-of-the-art method on the Hard-Impute subset of TrajImpute. We report the ADE metric. The unit for ADE is meters. The best results are bolded, and the second-best results are underlined.
Method ETH HOTEL UNV ZARA1 ZARA2 Average
[…]
Supervised-RealTraj 1.11 0.39 0.84 0.59 0.46 0.68
RealTraj 1.17 0.37 0.97 0.78 0.47 0.73
Additionally, we evaluate RealTraj’s robustness against missing data (i.e., miss-detection) using the TrajImpute [9]. Since the combination of the imputation method SAITS [13] and EqMotion achieved the best performance in the TrajImpute benchmark, we adopt SAITS as the imputation method in this evaluation. The results, presented in Tab. 3, show that RealTraj outperforms all comparison methods across all subsets. This includes significant improvements on the challenging HOTEL subset, where the imputation is especially difficult [9]. These findings confirm that the proposed model significantly enhances robustness against perception errors, aligning with our goal of addressing the first robustness limitation described in Sec. 1.
@@ 4.7 Ablation Study
Table 4: Ablation study of the proposed pretext tasks on the JRDB dataset. ’F’, ’P’, ’U’, and ’D’ represent forecasting, person ID reconstruction, unmasking, and denoising pretext tasks, respectively. Results are reported for both complete input and corrupted input scenarios with detection errors. The reported metric is ADE.
@@ F P U D Miss-Detection Error Localization
Effectiveness of Pretext Tasks. We first examine the effect of incorporating our proposed pretext tasks during pretraining on the JRDB, as shown in Tab. 4. In addition to using complete inputs, we conducted ablation with inputs containing detection errors to demonstrate the contribution of each pretext task to robustness. Our results show that all pretext tasks enhance trajectory forecasting performance. Specifically, the person identity reconstruction task improves ADE by $7.9\%$ (from $0.39$ to $0.36$). Both the unmasking and denoising tasks also contribute to performance gains, each enhancing ADE by $5.3\%$ (from $0.38$ to $0.36$). Joint training with all three pretext tasks and the main forecasting task improves ADE by $14\%$ (from $0.42$ to $0.36$) compared to pretraining with only the main forecasting task. Overall, the pretraining with our proposed pretext tasks and the main task improves the $54\%$ (from $0.78$ to $0.36$) on ADE compared to the model trained from scratch.
@@
Refer to caption (a) Effect of corruption ratio during training. Refer to caption (b) Effect of the number of synthetic trajectories.
@@
Effect of the Number of Synthetic Trajectories. Fig. 5 (b) presents the ADE transition when the number of pretraining samples is varied. Our framework shows improvement as the number of data increases.
@@
Figure 6: The visualization of predicted trajectories on the JRDB and JTA. The red circle indicates the detection of the target pedestrian in the last observed frame, and the gray circles represent past detections.
@@ 5 Qualitative Results
Fig. 6 visualizes the proposed method’s predicted results and the ground-truth future trajectories, highlighting the method’s accuracy in predicting future trajectories in complex scenes only using only past detections.
@@ 6 Limitation
Similar to conventional trajectory forecasting approaches, our method can only predict future trajectories for pedestrians detected in the last observed frame. Ideally, predicting future paths for all pedestrians observed in any past frame, even if occluded in the last observed frame, would enhance the applicability of our approach in complex scenes. Addressing this limitation presents an interesting direction for future work.
@@ 7 Conclusion
In this paper, we introduced RealTraj, a novel pedestrian trajectory forecasting framework designed to address three key limitations in conventional works. Our approach combines self-supervised pretraining on synthetic data with weakly-supervised fine-tuning on limited real-world data, reducing data collection requirements while enhancing robustness to perception errors. We proposed Det2TrajFormer, a model that leverages past detections to remain robust against tracking noise. By incorporating multiple pretext tasks during pretraining, we further improved the model’s robustness and forecasting accuracy using only detection inputs. Unlike existing methods, our framework fine-tunes only using ground-truth detections, reducing person ID annotation costs. We comprehensively validated its effectiveness against the limitations across multiple datasets, with experimental results demonstrating that our method outperforms state-of-the-art trajectory forecasting approaches.

### v3

@@ 4 Experiments
Refer to caption (a) Miss-Detection Refer to caption (b) Error Localization Refer to caption (c) Identity Switch Refer to caption (d) Combined Error
@@ 4.1 Datasets
Joint Track Auto Dataset (JTA): The JTA dataset [20] is a large-scale synthetic dataset comprising $256$ training sequences, $128$ validation sequences, and $128$ test sequences set in urban scenarios, generated using a video game engine.
ETH-UCY: The ETH and UCY datasets [56, 40] serve as established benchmarks for human trajectory forecasting, encompassing five scenes. We utilize the leave-one-out validation strategy as described in prior works [29, 47].
Stanford Drone Dataset (SDD): The SDD dataset [59] features over $11,000$ unique pedestrians across $20$ scenes, captured from a bird’s-eye view using a drone. Following the previous trajectory forecasting methods [47], we use the standard setup and train-test split.
TrajImpute: The TrajImpute dataset [13] simulates missing values within ETH-UCY, offering a framework for evaluating trajectory imputation and prediction models.
Following the standard practice, we predict $12$ future timesteps based on $9$ past timesteps for the JRDB and JTA datasets, and $12$ future timesteps based on $8$ past timesteps for the ETH-UCY, SDD, and TrajImpute datasets.
@@ 4.2 Evaluation Metrics
Average Displacement Error (ADE) and Final Displacement Error (FDE) are employed as evaluation metrics. ADE quantifies the average error over all predicted and ground-truth trajectory points, while FDE measures the displacement at the endpoints. Additionally, we report the minimum ADE and FDE (minADEN, minFDEN) over $N=20$ trajectories per pedestrian for multi-future prediction setting.
@@ 4.3 Implementation Details
We use the ORCA crowd simulator [70, 58] to generate 2K synthetic trajectories, as shown in Fig. 3. We train our model using the Adam optimizer [37], starting with a learning rate of $1\times 10^{-4}$, which is reduced by a factor of $0.1$ after completing $80\%$ of the total $200$ epochs for pretraining and $20$ epochs for fine-tuning. We set $\lambda=10$ for balancing the loss weight in Eq. 6. See the supplementary material for more details (e.g., training hyperparameters).
4.4 Robustness Evaluation
The robustness of the proposed method against perception errors (detection and tracking errors) is compared with that of the transformer-based (Social-Trans [61]) and graph-based (EqMotion [79]) models, which are the well-known state-of-the-art methods across multiple benchmarks (c.f., Tab. 3). Although Social-Trans is designed to process multiple modalities, we used only trajectory data as input in all experiments to ensure fair comparisons. We synthetically generated two types of detection errors—miss-detections and localization errors—along with one type of tracking error, specifically identity switches, and combined these errors for evaluation. Miss-detections were simulated by setting past coordinates to zero for our model and linearly interpolated for comparison models, while localization errors were introduced by adding Gaussian noise to the inputs. Identity switches were simulated by swapping identities with nearby pedestrians within a 5-meter radius. Both detection and tracking errors were introduced for combined errors.
Table 1: Comparison with the current state-of-the-art method on the Hard-Impute subset of TrajImpute. We report the minADE20 metric. The unit for minADE20 is meters. The best results are bolded, and the second-best results are underlined. Note that our model takes detections as inputs, while other approaches take the trajectory as inputs.
Method ETH HOTEL UNV ZARA1 ZARA2 Avg.
GPGraph [3] 0.92 1.89 0.53 0.58 0.36 0.86
GraphTern[2] 0.78 1.68 0.50 0.96 0.37 0.86
LBEBM-ET [4] 0.85 3.31 0.64 0.37 0.27 1.09
SGCN-ET [4] 1.07 3.21 0.77 0.61 0.41 1.21
TUTR [64] 1.12 3.36 0.59 0.50 0.33 1.18
EqMotion [79] 0.47 0.72 0.39 0.28 0.37 0.45
RealTraj (Ours) 0.48 0.20 0.37 0.28 0.21 0.30
Additionally, we evaluate RealTraj’s robustness against missing data (i.e., miss-detection) using the TrajImpute [13]. Since the combination of the imputation method SAITS [18] and EqMotion achieved the best performance in the TrajImpute benchmark, we adopt SAITS as the imputation method in this evaluation. The results, presented in Tab. 1, show that RealTraj outperforms all comparison methods across most subsets. This includes significant improvements on the challenging HOTEL subset, where the imputation is especially difficult [13]. Additional result under the momentary trajectory prediction and observation length shift settings is provided in the supplementary materials. These findings confirm that the proposed model significantly enhances robustness against perception errors, aligning with our goal of addressing the first robustness limitation described in Sec. 1.
Table 2: Comparisons on JRDB, JTA, ETH-UCY, and SDD under few-shot evaluation setting. The ADE metric is reported with different percentages of labeled real data during fine-tuning. The S and WS in the Setting column indicate that the models are trained in a fully-supervised and weakly-supervised manner, respectively. The T and D in the Input column denote that the models take trajectories and detections as inputs, respectively. The column (w/ Syn.) indicates whether the models are pretrained on the synthetic trajectory data. The unit for ADE is meters for JRDB, JTA, and ETH-UCY, and pixels for SDD. The best results are bolded, and the second-best results are underlined.
(a) JRDB
[…]
S T EqMotion [79] 0.45 0.48 0.43 0.60 0.43
0.46 0.45 0.42 0.41 0.41
Social-Trans [61] 0.68 0.48 0.47 0.43 0.43
0.61 0.45 0.43 0.42 0.41
0.92 0.88 0.88 0.87 0.86
D RealTraj (Ours) 0.43 0.41 0.40 0.38 0.38
1.11 0.89 0.88 0.87 0.86
WS D RealTraj (Ours) 0.43 0.41 0.40 0.39 0.38
(b) JTA
[…]
S T EqMotion [79] 1.76 1.68 1.55 1.55 1.55
1.75 1.55 1.53 1.46 1.43
Social-Trans [61] 4.80 3.33 2.91 2.78 2.78
2.48 1.63 1.49 1.39 1.29
4.83 4.74 4.67 1.43 1.37
D RealTraj (Ours) 1.38 1.25 1.22 1.20 1.07
4.98 4.93 4.91 4.88 4.86
WS D RealTraj (Ours) 2.13 1.98 1.68 1.46 1.23
(c) ETH-UCY
[…]
S T EqMotion [79] 0.70 0.56 0.53 0.53 0.53
0.81 0.60 0.60 0.61 0.56
Social-Trans [61] 1.05 0.62 0.55 0.52 0.53
0.57 0.56 0.52 0.52 0.50
0.81 0.72 0.60 0.61 0.57
D RealTraj (Ours) 0.56 0.55 0.53 0.52 0.52
1.37 1.13 0.81 0.67 0.57
WS D RealTraj (Ours) 0.56 0.56 0.56 0.53 0.53
(d) SDD
[…]
S T EqMotion [79] 24.7 20.0 18.5 19.0 17.6
25.8 20.9 19.3 19.2 19.8
Social-Trans [61] 45.0 32.7 20.5 21.8 19.0
31.5 18.1 17.9 17.4 17.3
67.4 41.8 38.3 36.3 18.7
D RealTraj (Ours) 22.3 20.2 19.2 18.5 17.0
70.6 54.4 41.2 38.7 19.5
WS D RealTraj (Ours) 23.5 20.3 19.7 19.0 17.9
Table 3: Comparisons with state-of-the-art methods on the JTA, JRDB, ETH-UCY, and SDD under fully-supervised setting. The S and WS in the Setting column indicate that the models are trained in a fully-supervised and weakly-supervised manner, respectively. The T and D in the Input column denote that the models take trajectories and detections as inputs, respectively. The unit for ADE and FDE is meters for JRDB, JTA, and ETH-UCY, and pixels for SDD. The best results are bolded, and the second-best results are underlined. †: We employ Social-Transmotion with trajectory inputs for a fair comparison. For the multi-future prediction setting, we extend it to generate multiple future trajectories using the same approach as ours.
Setting Input Method JRDB JTA ETH-UCY SDD
ADE FDE ADE FDE ADE FDE minADE20 minFDE20 ADE FDE minADE20 minFDE20
S T Social-LSTM [1] 0.47 0.95 1.21 2.54 0.72 1.54 - - - - - -
Social-GAN-ind [29] 0.50 0.99 1.66 3.76 0.74 1.54 0.58 1.18 - - 27.2 41.4
Trajectron++ [63] 0.40 0.78 1.18 2.53 0.53 1.11 0.21 0.41 - - - -
Transformer [27] 0.56 1.10 1.56 3.54 0.54 1.17 0.31 0.55 - - - -
Directional-LSTM [39] 0.45 0.87 1.37 3.06 - - - - - - - -
Autobots [26] 0.39 0.80 1.20 2.70 0.52 1.05 - - - - - -
TUTR [64] - - - - 0.53 1.12 0.21 0.36 17.4 35.0 7.76 12.7
EigenTrajectory [4] - - - - 0.51 1.11 0.22 0.37 20.7 41.9 8.12 13.1
FlowChain [45] - - - - - - 0.29 0.52 - - 9.93 17.17
EqMotion [79] 0.42 0.78 1.13 2.39 0.49 1.03 0.21 0.35 16.8 33.7 8.45 14.1
Social-Trans [61] 0.40 0.77 0.99 1.98 0.51 1.03 0.21 0.41 18.0 38.7 7.21 14.3
MGF [10] - - - - - - 0.21 0.34 - - 7.74 12.1
D RealTraj (Ours) 0.35 0.67 0.92 1.84 0.51 1.03 0.23 0.42 16.0 31.8 7.18 12.4
WS D RealTraj (Ours) 0.36 0.71 1.04 2.11 0.52 1.07 0.26 0.43 16.0 32.0 8.21 13.8
@@ 4.5 Few-shot Evaluation
Next, we demonstrate the effectiveness of RealTraj with limited real-world labeled data. We randomly selected subsets of $0.1\%$, $0.5\%$, $1\%$, $2\%$, and $5\%$ of the datasets, fine-tuned the model on each subset, and evaluated it on the test set on JRDB, JTA, ETH-UCY, and SDD. In this experiment, since no prior work focuses on trajectory prediction with a few-shot setting, we compare our model against the existing approaches, EqMotion [79] and Social-Trans [61]. We prepare the comparison models that are pretrained on the synthetic trajectory data, the same as ours, for a fair comparison.
As shown in Tab. 2, our method consistently outperforms these prior methods on JRDB. In the most challenging scenario with only $0.1\%$ of annotations, weakly-supervised RealTraj (last row) achieves improvements of $4.4\%$, $20\%$, and $4.9\%$ on JRDB, ETH-UCY, and SDD, respectively, compared to the best-performing previous supervised model. In other data settings, RealTraj achieves results comparable to supervised methods despite the detection inputs and requiring only weak detection annotations during fine-tuning. This demonstrates that our model overcomes the second limitation of real-world data collection cost mentioned in  Sec. 1. Furthermore, the weakly-supervised RealTraj (last row) achieved the performance on par with the fully-supervised RealTraj on JRDB and SDD datasets, verifying that our approach reduces the person ID annotation cost mentioned in  Sec. 1.
@@ 4.6 Comparison with State-of-the-art
We further compare RealTraj with a diverse range of existing approaches under the standard 100% labeled setting, as summarized in Tab. 3. Although achieving state-of-the-art performance in a fully supervised setting with clean trajectory inputs is not the primary focus of our paper, the results show that our weakly supervised framework performs comparably to fully supervised approaches using only detection inputs on the various benchmarks. Additionally, our fully supervised approach surpasses all other methods on the JRDB, JTA, and SDD datasets in terms of ADE and remains competitive on the ETH-UCY dataset.
@@ 4.7 Ablation Study
Table 4: Ablation study of the proposed pretext tasks on the JRDB dataset. ’F’, ’P’, ’U’, and ’D’ represent forecasting, person ID reconstruction, unmasking, and denoising pretext tasks, respectively. The first row denotes the results without pretraining phase. Results are reported for both complete input and corrupted input scenarios with detection errors. The reported metric is ADE.
@@ F P U D Miss-Detection Error Localization
Effectiveness of Pretext Tasks. We first examine the effect of incorporating our proposed pretext tasks during pretraining on the JRDB, as shown in Tab. 4. In addition to using complete inputs, we conducted ablation with inputs containing detection errors to demonstrate the contribution of each pretext task to robustness. Our results show that all pretext tasks enhance trajectory forecasting performance. First, the performance significantly degrades when removing the pretraining stage even with the complete input, verifying the effectiveness of pretraining for enabling the trajectory prediction with detections. Moreover, the person identity reconstruction task improves ADE by $7.9\%$ (from $0.39$ to $0.36$). Both the unmasking and denoising tasks also contribute to performance gains, each enhancing ADE by $5.3\%$ (from $0.38$ to $0.36$). Joint training with all three pretext tasks and the main forecasting task improves ADE by $14\%$ (from $0.42$ to $0.36$).
@@
Refer to caption (a) Effect of corruption ratio during training. Refer to caption (b) Effect of the number of synthetic sequences.
@@
Effect of the Number of Synthetic Sequences. Fig. 5 (b) presents the ADE transition when the number of pretraining samples is varied. Our framework shows improvement as the amount of data increases.
@@
Figure 6: The visualization of predicted trajectories on the JRDB and JTA. The red circle is the detection of the target pedestrian in the last observed frame, and the gray circles are past detections.
@@ 5 Qualitative Results
Fig. 6 visualizes the proposed method’s predicted results and the ground-truth future trajectories, highlighting the method’s accuracy in predicting future trajectories in complex scenes only using past detections as inputs.
@@ 6 Limitation
Similar to conventional trajectory forecasting models, our method can only predict future trajectories for pedestrians detected in the last observed frame. Ideally, predicting future paths for all pedestrians observed in any past frame, even if occluded in the last observed frame, would enhance the applicability in complex scenes. Addressing this limitation presents an interesting direction for future work.
@@ 7 Conclusion
In this paper, we introduced RealTraj, a novel pedestrian trajectory forecasting framework designed to address three key limitations in conventional works. Our approach combines self-supervised pretraining on synthetic data with weakly-supervised fine-tuning on limited real-world data, reducing data collection requirements while enhancing robustness to perception errors. We proposed Det2TrajFormer, a model that leverages past detections to remain robust against tracking noise. By incorporating multiple pretext tasks during pretraining, we further improved the model’s robustness and forecasting accuracy using only detection inputs. Unlike existing methods, our framework fine-tunes only using ground-truth detections, reducing person ID annotation costs. In our experiments, we thoroughly validated that our approach effectively overcomes the key limitations of existing trajectory prediction models, particularly when applied to real-world scenarios.
Supplementary Material
8 Methodological Comparison of Robust Trajectory Forecasting against Perception Errors
To offer a clear and concise comparison between prior works addressing perception errors and our RealTraj approach, we have summarized the key differences in Tab. 6. Unlike previous methods that typically employ architectures and objectives tailored to either detection or tracking errors, RealTraj aims to develop a robust model capable of addressing both detection and tracking errors.
Table 6: Comparative overview of methodological features for robust trajectory forecasting against perception errors.
Methods Detection Errors Tracking Error
Miss-Detection Error Localization Identity Switch
[23, 82, 86, 65, 80, 67, 42, 52] - -
N/A - -
[84, 73, 72, 88, 21] - -
RealTraj (Ours)
Refer to caption (a) Miss-Detection Refer to caption (b) Error Localization Refer to caption (c) Identity Switch Refer to caption (d) Combined Error
Figure 7: Comparison of model robustness against various error types on the JTA dataset, including detection errors (miss-detections and localization errors), tracking errors (identity switches), and their combined impact.
9 Implementation Details
As noted, we use the ORCA crowd simulator [70], implemented by [58], to generate synthetic trajectories. Up to $40$ pedestrians are placed in a $15\text{m}\times 15\text{m}$ environment with up to $20$ static primitive obstacles. We generate $2000$ sequences for pretraining.
During the first $100$ epochs of pretraining, the loss weights $\alpha$, $\beta$, and $\gamma$ are set to $1$, 00, and 00, respectively; for the final $100$ epochs, these weights are set to 00, $100$, and $0.1$. A corruption ratio of $30\%$ is applied. Hyperparameters were selected through a coarse-to-fine grid search and step-by-step tuning. The model is trained with a batch size of $16$ on a single NVIDIA TITAN RTX GPU.
The model configuration for Det2TrajFormer consists of nine layers and four attention heads, with a model dimension of $d=128$. The person ID embedding dimension $d_{I}$ is set to $64$. During training, we corrupt past coordinates with a corruption probability $p_{c}$ of $30\%$. This corruption is introduced by either simulating miss-detections or adding localization errors. Miss-detections are emulated by setting the coordinate values to zero, while localization errors are simulated by adding Gaussian noise $\epsilon\sim\mathcal{N}(0,\sigma^{2})$, where the standard deviation $\sigma$ is set to $0.5$.
10 Additional Experimental Results
Robustness Evaluation We assess the impact of detection and tracking errors introduced in the JTA dataset, as shown in Fig. 7. The proposed method exhibits only minor performance degradation, highlighting its robustness to detection errors. Notably, RealTraj consistently outperforms all baselines across all error ratios in the JTA dataset.
Momentary Trajectory Prediction. We evaluate our model in a momentary trajectory prediction setting, where only two frames serve as input, and the model predicts future trajectories for the next 12 timesteps. The results in Tab. 7 show that our model, which integrates three key limitations into a unified framework, achieves performance comparable to specialized momentary prediction approaches.
Observation Length Shift. To examine the model’s adaptability to observation length shift, we test it on observation lengths different from those used during training. As shown in Fig. 8, our model consistently outperforms all baselines across varying observation lengths.
Few-shot Evaluation. In addition to the results presented in Table 1 of the main paper, we further investigated the capabilities of our model using limited real-world labeled data by randomly selecting subsets containing $5\%$, $10\%$, $20\%$, $40\%$, and $60\%$ of the dataset. The model was fine-tuned on each subset and subsequently evaluated on the test set. As shown in Fig. 9, our RealTraj consistently outperforms both EqMotion and Social-Trans, benefiting from pre-trained knowledge.
Comparison of fine-tune and scratch model. We conducted an ablation study to evaluate whether the model benefits from pre-trained knowledge. We compared the performance of a fine-tuned model trained from a pretrained model to models trained from scratch. As shown in  Tab. 8, the fine-tuned model consistently outperformed the scratch models by a significant margin. This suggests that pre-training with our proposed pretexts effectively enhances trajectory forecasting performance.
Table 7: Comparison with the current state-of-the-art methods on ETH-UCY and SDD for momentary trajectory prediction. The best results are bolded, while the second-best results are underlined.
Method ETH-UCY SDD
minADE20 minFDE20 minADE20 minFDE20
MOE [67] 0.20 0.41 8.40 16.08
DTO [52] 0.23 0.46 8.32 11.56
BCDiff [42] 0.19 0.39 8.93 16.92
Supervised-RealTraj 0.24 0.39 8.05 13.32
Figure 8: Comparisons with the current state-of-the-art method on JRDB with different observation lengths. We report the ADE metric.The unit for ADE is meters.
Figure 9: Comparisons with the current state-of-the-art method on JRDB under few-shot evaluation. We summarize the ADE metric with different percentages of labeled data. The unit for ADE is meters.
Table 8: Comparison of fine-tune and scratch model on JRDB, JTA, ETH-UCY, and SDD. We report the ADE metric. The unit for ADE is meters.
Method JRDB JTA ETH-UCY SDD
Scratch 0.78 1.15 1.05 17.74
Fine-tune (RealTraj) 0.36 1.04 0.52 16.02
Table 9: Performance results for different numbers of encoder layers in Det2TrajFormer. An encoder depth of 9 layers achieves the best balance between performance and computational efficiency.
L ADE FDE
3 0.38 0.76
6 0.37 0.71
9 0.36 0.69
12 0.36 0.69
11 Effect of Number of Encoder Layers
As shown in Tab. 9, a relatively deep encoder is crucial for optimal performance. Increasing the number of encoder layers from $3$ to $9$ results in a $5.3\%$ improvement in ADE. However, adding more layers beyond this point does not yield significant gains. Therefore, we use $9$ encoder layers as our default setting, offering a favorable balance between efficiency and performance.